# Garf: Learning Generalizable 3D Reassembly for Real-World Fractures

### v1

@@ B.3 Preliminaries on Riemannian Flow Matching
Instead of simulating discrete noise addition steps, flow matching (FM) learns a probability density path $p_{t}$, which progressively transforms a noise distribution $p_{t=0}$ to the data distribution $p_{t=1}$, with a time variable $t\in[0,1]$. As a simulation-free method aiming to learn continuous normalizing flow (CNF), FM models a probability density path $p_{t}$, which progressively transforms a noise distribution $p_{t=0}$ to the data distribution $p_{t=1}$, with a time variable $t\in[0,1]$. Inspired by learning assembly by breaking, the rigid motion of the fragments corresponds to the geodesic on the Lie group $\mathrm{SE}(3)$, which is a differentiable Riemannian manifold. Inspired by previous works [4, 58, 11], FM can be extended to $\mathrm{SE}(3)$ manifold to learn the rigid assembly process.
@@
Table III: Results on Vanilla Breaking Bad [42] Dataset.
@@
degree ×$\times 10^{-2}$$\times 10^{-3}$
@@
Global [22] 80.70 15.10 24.60 14.60
LSTM [53] 84.20 16.20 22.70 15.80
DGL [60] 79.40 15.00 31.00 14.30
SE(3)-Equiv [54] 79.30 16.90 8.41 28.50
DiffAssemble [41] 73.30 14.80 27.50 -
PHFormer [7] 26.10 9.30 50.70 9.60
Jigsaw [30] 42.30 10.70 57.30 13.30
PF++ [51] 38.10 8.04 70.60 6.03
@@ C Additional Results and Analyses
C.1 Quantitative Results on Vanilla Breaking Bad
@@
Given that all our previous experiments were conducted on the volume-constrained version of the Breaking Bad dataset [42], we here provide additional quantitative results on the non-volume-constrained version to align with the settings of previous methods. The results, shown in Table III, demonstrate that our Garf-mini model still significantly outperforms the previous state-of-the-art method, PF++ [51], by a large margin. This performance is consistent across both the everyday and artifact subsets, showcasing the model’s robust generalization ability.
@@
We also present the results of FragmentDiff [57] on their custom Breaking Bad dataset in Table IV. FragmentDiff claims to remove tiny pieces, but it is unclear whether this applies only to their training setting or also to evaluation. Unfortunately, since they did not open source their code or provide their preprocessed data, we are unable to directly compare all other methods with FragmentDiff. Additionally, they did not adhere to the common settings used by other methods, which limit the number of pieces from 2 to 20, making direct comparisons on their provided metrics impossible. However, its significant performance drop from the Everyday subset to the Artifact subset suggests that Garf surpasses FragmentDiff in generalization capability.
C.2 Quantitative Results of Finetuning on the Fractura Synthetic Dataset
@@
After finetuning Garf on the Fractura synthetic dataset, we report the per-category performance on the bone and eggshell categories, as shown in Table V. The results demonstrate that finetuning the FM model in Garf significantly improves performance on these two unseen categories, showing the effectiveness of our finetuning techniques and the generalizability of our pretraining strategy.
C.3 Additional Qualitative Comparison on the Fractura and Breaking Bad Dataset
@@
Figures IIIII and IV demonstrate more qualitative comparison on the Fractura and Breaking Bad Dataset, where our Garf shows superior performance than the other previous SOTA methods.
@@
Table IV: FragmentDiff [57] Results on Their Custom Breaking Bad Dataset.
@@
degree ×$\times 10^{-2}$
FragmentDiff [57] Everyday 13.68 7.41 90.20
@@
Table V: Quantitative Per-category Results on the Fractura (Synthetic Fracture).
@@
degree ×$\times 10^{-2}$$\times 10^{-3}$
@@
Figure II: Qualitative Results on the Fractura Synthetic Dataset.
@@
Figure IV: Qualitative Results on the Breaking Bad Dataset Artifact Subset.

### v2

@@ B.3 Preliminaries on Riemannian Flow Matching
Instead of simulating discrete noise addition steps, flow matching (FM) learns a probability density path $p_{t}$, which progressively transforms a noise distribution $p_{t=0}$ to the data distribution $p_{t=1}$, with a time variable $t\in[0,1]$. As a simulation-free method aiming to learn continuous normalizing flow (CNF), FM models a probability density path $p_{t}$, which progressively transforms a noise distribution $p_{t=0}$ to the data distribution $p_{t=1}$, with a time variable $t\in[0,1]$. Inspired by learning assembly by breaking, the rigid motion of the fragments corresponds to the geodesic on the Lie group $\mathrm{SE}(3)$, which is a differentiable Riemannian manifold. Inspired by previous works [bose2023se, yim2023fast, geffnerproteina], FM can be extended to $\mathrm{SE}(3)$ manifold to learn the rigid assembly process.
@@
B.4 More attention on large fragments
Garf provides tailored designs to place more attention on large fragments. We observed that: (i) large fragments are easier to assemble; (ii) tiny fragments sometimes lead to unstable training. Driven by these insights, we apply weighted sampling based on the surface area of fragments and modify the self-attention module to allow more attention on large fragments.
Figure II: Self-attention comparison between Garf (left) and PuzzleFusion++ [wang2024puzzlefusion++] (right).
Table III: Results on Vanilla Breaking Bad [sellan2022breaking] Dataset.
@@ B.4 More attention on large fragments
degree $\times 10^{-2}$ % $\times 10^{-3}$
@@
Global [li2020learning] 80.70 15.10 24.60 14.60
LSTM [wu2020pq] 84.20 16.20 22.70 15.80
DGL [zhan2020generative] 79.40 15.00 31.00 14.30
SE(3)-Equiv [wu2023leveraging] 79.30 16.90 8.41 28.50
DiffAssemble [scarpellini2024diffassemble] 73.30 14.80 27.50 -
PHFormer [cui2024phformer] 26.10 9.30 50.70 9.60
Jigsaw [lu2024jigsaw] 42.30 10.70 57.30 13.30
PF++ [wang2024puzzlefusion++] 38.10 8.04 70.60 6.03
@@
Table IV: Ablation Study on Our Designs of FM.
10.24 1.95 89.08
8.02 1.63 93.78
7.63 1.60 94.02
6.68 1.34 94.77
Table V: Ablation Study on Sample Steps.
1 12.52 3.18 86.88 2.14 38.26
One-Step + 1 9.79 2.46 91.31 1.42 45.76
5 8.25 1.92 93.70 0.53 57.32
One-Step + 5 7.15 1.66 94.43 0.46 76.23
20 7.63 1.60 94.02 0.35 185.05
One-step + 20 6.68 1.34 94.77 0.25 190.77
50 7.50 1.54 94.01 0.32 408.40
Table VI: Ablation Study on the Different Anchor Initialization.
Largest Anchor 6.10 1.22 95.33 0.22
Random Anchor 6.09 1.30 95.20 0.29
Anchor-Free 9.09 2.13 93.23 0.91
Table VII: Comparison Between Diffusion and Our FM Models.
Dataset Methods RMSE(R) $\downarrow$ RMSE(T) $\downarrow$ PA $\uparrow$
Everyday Diffusion 7.45 1.47 94.30
$\mathrm{SE}(3)$ Diffusion N/A N/A N/A
Diffusion w/ One-Step 7.51 1.47 94.27
Vanilla FM 10.24 1.95 89.08
Garf-mini 6.68 1.34 94.77
Fractura Diffusion $32.38$ $7.90$ $71.73$
Garf-mini 27.88 6.79 76.25
@@ C Additional Results and Analyses
C.1 Ablation on Design Choices in Flow Matching
We conduct an ablation study to evaluate the impact of design choices in our FM module. As shown in Table IV, vanilla FM, trained with spherical linear interpolation (slerp) to approximate valid rotations in the forward process [guo2025assembleflow], achieves 89.08 PA, already surpassing previous methods [wang2024puzzlefusion++, lu2024jigsaw]. Incorporating the $\mathrm{SE}(3)$ representation further improves performance by pre-modeling the manifold distribution and better capturing distribution shifts during assembly. Multi-anchor training strategy further enhances results, while one-step pre-assembly significantly boosts performance by providing a more reasonable initial pose distribution, leading to the best overall outcomes.
C.2 Ablation on Sample Steps
Table V shows the effect of varying sampling steps in our framework. Surprisingly, even with just 5 steps, FM achieves 93.70% PA, highlighting its effectiveness in modeling global probabilistic paths. Additionally, our first-session initialization provides a more reasonable initial pose, further improving assembly quality while adding minimal computational overhead.
C.3 Ablation on Anchor Fragment
Similar to PF++ [wang2024puzzlefusion++], we use the largest fragment as the anchor fragment at inference. We compare the performance of using the largest fragment, a randomly selected fragment, and no anchor fragment. As shown in Table VI, using a random fragment as the anchor fragment has almost no negative effect on the model. Only anchor-free initialization leads to a slight performance drop.
C.4 Comparison with Diffusion Models
Table VII compares our FM module with diffusion models. While diffusion, when paired with fracture-aware pretraining, achieves competitive performance, directly applying vanilla FM yields lower results (89.08 PA), emphasizing the importance of our subsequent design choices. A key limitation of diffusion models is their handling of $\mathrm{SO}(3)$ rotation, which cannot be naturally incorporated into the reverse process. Existing methods, such as score prediction [yim2023se], aim to maintain rotation validity but fall outside our current scope. Additionally, diffusion models rely on multi-step denoising without explicitly modeling the global probabilistic path, rendering one-step pre-assembly ineffective. Furthermore, on Fractura, diffusion models exhibit weaker generalization to unseen objects compared to Garf-mini.
C.5 Quantitative Results on Vanilla Breaking Bad
@@
Given that all our previous experiments were conducted on the volume-constrained version of the Breaking Bad dataset [sellan2022breaking], we here provide additional quantitative results on the non-volume-constrained version to align with the settings of previous methods. The results, shown in Table III, demonstrate that our Garf-mini model still significantly outperforms the previous state-of-the-art method, PF++ [wang2024puzzlefusion++], by a large margin. This performance is consistent across both the everyday and artifact subsets, showcasing the model’s robust generalization ability.
@@
We also present the results of FragmentDiff [xu2024fragmentdiff] on their custom Breaking Bad dataset in Table VIII. FragmentDiff claims to remove tiny pieces, but it is unclear whether this applies only to their training setting or also to evaluation. Unfortunately, since they did not open source their code or provide their preprocessed data, we are unable to directly compare all other methods with FragmentDiff. Additionally, they did not adhere to the common settings used by other methods, which limit the number of pieces from 2 to 20, making direct comparisons on their provided metrics impossible. However, its significant performance drop from the Everyday subset to the Artifact subset suggests that Garf surpasses FragmentDiff in generalization capability.
C.6 Quantitative Results of Finetuning on the Fractura Synthetic Dataset
@@
After finetuning Garf on the Fractura synthetic dataset, we report the per-category performance on the bone and eggshell categories, as shown in Table IX. The results demonstrate that finetuning the FM model in Garf significantly improves performance on these two unseen categories, showing the effectiveness of our finetuning techniques and the generalizability of our pretraining strategy.
C.7 Additional Qualitative Comparison on the Fractura and Breaking Bad Dataset
@@
Figures IIIIV and V demonstrate more qualitative comparison on the Fractura and Breaking Bad Dataset, where our Garf shows superior performance than the other previous SOTA methods.
@@
Table VIII: FragmentDiff [xu2024fragmentdiff] Results on Their Custom Breaking Bad Dataset.
@@
degree $\times 10^{-2}$ %
FragmentDiff [xu2024fragmentdiff] Everyday 13.68 7.41 90.20
@@
Table IX: Quantitative Per-category Results on the Fractura (Synthetic Fracture).
@@
degree $\times 10^{-2}$ % $\times 10^{-3}$
@@
Figure IV: Qualitative Results on the Fractura Synthetic Dataset.
Figure V: Qualitative Results on the Breaking Bad Dataset Artifact Subset.